OpenStudy (anonymous):

which is an equation in point-slope form for the given point and slope? Point: ( 2, 6) ; Slope : -3/4

OpenStudy (anonymous):

y-y1=m(x-x1) y-6=-3/4(x-2) y-6=-3/4x+6/4

OpenStudy (anonymous):

4y-24=-3x+6 4y=-3x+6+24 4y=-3x+30

OpenStudy (anonymous):

3x+4y=30